The larger of two numbers is 11 more than twice the smaller. The sum of the numbers is 1 less than seven times smaller. Find the numbers.
Answer by gonzo(654)

L = larger number
S = smaller number
-----
L = 2*S + 11
L + S = 7 * S - 1
-----
substituting for L in the second equation by using the equality in the first equation, we get
(2*S + 11) + S = 7*S - 1
removing parentheses, we get
2*S + 11 + S = 7*S - 1
combining like terms, we get
3*S + 11 = 7*S - 1
subtracting 3*S from both sides of the equation, we get
11 = 7*S - 1 - 3*S
adding 1 to both sides of the equation, we get
11 + 1 = 7*S - 3*S
combining like terms, we get
12 = 4*S
dividing both side of equation by 4, we get
3 = S
which is the same as
S = 3
-----
substituting for S in the first equation, we get
first equation:
L = 2*S + 11
L = 2*3 + 11
L = 6 + 11
L = 17
-----
substituting for S and L in the second equation, we get
L + S = 7*S - 1
17 + 3 = 7*3 - 1
20 = 21 - 1
20 = 20
answer is correct
-----
answer is:
L = 17
S = 3